what is sum of the first 20 terms if the firs term is 73 and the last term is -101?
Answer by jim_thompson5910(35256)

Sn = (n*(a1 + an))/2

S20 = (20*(73 + -101))/2

S20 = (20*(-28))/2

S20 = (-560)/2

S20 = -280

So the answer is -280